# GearTrack: Automating 6D Pose Estimation

### v1

@@ 1 Introduction
Ultimately, we found that SAM2 (Ravi et al., 2024) provided superior results. By simply clicking on the location of the object in the test image, the algorithm automatically generates a segmentation and a ‘pure’ image based on the user-specified object. For future runs, only the CAD model is needed, and the system leverages LightGlue (Lindenberger et al., 2023) to match feature points between the ‘pure’ image and the first frame of the live video or image. These matched points serve as location prompts for SAM2, enabling fully automated object tracking and segmentation without any additional user input. This approach not only addresses the mask annotation challenge but also facilitates a fully automated workflow for object pose estimation.
Figure 1: System workflow The image shows the implementation process of the entire system. In our system, the object’s positional information within the real image is initially obtained either through user clicks or by employing LightGlue to perform feature matching between the pure image and the real image, thus providing a positional prompt. This information is then transmitted to SAM2. If a pure image of the object does not exist, one is generated and stored in memory. Simultaneously, a mask matrix for the object is created. Subsequently, by integrating the generated mask, the object’s CAD model, and the real frame, FoundationPose is ultimately utilized to perform 6D pose estimation.
As shown in Figure 1, we only need to provide a CAD model and click on the target we want to recognize during the first use or provide a pure image of the target. The system can then automatically complete the entire 6D pose estimation without any other operations.
@@ 3.1 Integration and Testing of Methods
3.1.1 Initial Target Identification via Manual Selection or Pure Image Input
The system requires the user to provide a CAD model of the object to be recognized. During the first run, the user can either manually select the target object by clicking on it or provide a pure image of the object.
@@ 3.1.2 Segmentation Prompt Generation Based on User Input or Feature Matching
Once the target object is identified via manual selection or pure image input(an image of the object with a pure white background), the system generates prompt information used by SAM2 for segmentation. SAM2, trained on the SA-V dataset(Ravi et al., 2024), utilizes a Transformer-based encoder with ViT(Ravi et al., 2024) to extract image features and applies a Memory Attention mechanism to handle dependencies in video data. For mask, point, and box-type prompts, the encoder encodes these prompts, and the resulting representations, along with image features, are fed into a mask decoder to generate the mask. Thus, the object’s location, determined either by the user’s click or through feature point matching via LightGlue, serves as a point-type prompt for SAM2’s mask decoder, which then generates the final object mask.
@@ 3.2 Handling Tracking Loss in FoundationPose
FoundationPose relies on positional information from the previous frame for pose estimation, which can result in tracking loss if the object moves too quickly. To address this, SAM2 performs real-time object segmentation for each frame. We calculate the distance between the centroid of the mask and the center point of the pose estimate to detect tracking loss. In Eq.1, this distance is measured using Robust Lorentzian centroid Distance with Lorentzian error function and square of L2 norm.
@@ 3.3 Addressing Long-Term Object Loss: A Memory Mechanism
To handle long-term object loss, we implemented a memory-like mechanism. As defined in Eq.3, if the object is lost for an extended period (typically 10 seconds), both FoundationPose and SAM2 may fail to track it.
@@
In such cases, we reload the pure image of the object saved during the initial SAM2 run and perform feature point matching on the current frames using LightGlue. If the object reappears, a location prompt is generated, which is passed to SAM2 for re-segmentation. The segmentation result is then fed into FoundationPose for re-registration, allowing the system to resume tracking the object.

### v2

# SuperPose: Improved 6D Pose Estimation with Robust Tracking and Mask-Free Initialization

@@ 1 Introduction
Ultimately, we found that SAM2 (Ravi et al., 2024) provided superior results. By simply clicking on the location of the object in the test image, the algorithm automatically generates a segmentation and a segmented image based on the user-specified object. For future runs, only the CAD model is needed, and the system leverages LightGlue (Lindenberger et al., 2023) to match feature points between the segmented image and the first frame of the live video or image. These matched points serve as location prompts for SAM2, enabling fully automated object tracking and segmentation without any additional user input. This approach not only addresses the mask annotation challenge but also facilitates a fully automated workflow for object pose estimation.
Figure 1: System workflow The image shows the implementation process of the entire system. In our system, the object’s positional information within the real image is initially obtained either through user clicks or by employing LightGlue to perform feature matching between the segmented image and the real image, thus providing a positional prompt. This information is then transmitted to SAM2. If a segmented image of the object does not exist, one is generated and stored in memory. Simultaneously, a mask matrix for the object is created. Subsequently, by integrating the generated mask, the object’s CAD model, and the real frame, FoundationPose is ultimately utilized to perform 6D pose estimation.
As shown in Figure 1, we only need to provide a CAD model and click on the target we want to recognize during the first use or provide a segmented image of the target. The system can then automatically complete the entire 6D pose estimation without any other operations.
@@ 3.1 Integration and Testing of Methods
Figure 2: The initialization via manual selection or segmented image input The image illustrates the initialization process of our system. In the first method, users simply click on the frame, which sends a prompt to SAM2. This generates a mask that is passed to FoundationPose to produce a pose estimation and a segmented image, which is then stored for the memory mechanism. In the second method, users provide a segmented image to LightGlue, which generates a prompt for SAM2. SAM2 then produces a mask, which is used by FoundationPose to generate the estimation.
3.1.1 Initial Target Identification via Manual Selection or segmented Image Input
The system requires the user to provide a CAD model of the object to be recognized. During the first run, the user can either manually select the target object by clicking on it or provide a segmented image of the object as demonstrated in Figure 2.
@@ 3.1.2 Segmentation Prompt Generation Based on User Input or Feature Matching
Once the target object is identified via manual selection or segmented image input(an image of the object with a pure white background), the system generates prompt information used by SAM2 for segmentation. SAM2, trained on the SA-V dataset(Ravi et al., 2024), utilizes a Transformer-based encoder with ViT(Ravi et al., 2024) to extract image features and applies a Memory Attention mechanism to handle dependencies in video data. For mask, point, and box-type prompts, the encoder encodes these prompts, and the resulting representations, along with image features, are fed into a mask decoder to generate the mask. Thus, the object’s location, determined either by the user’s click or through feature point matching via LightGlue, serves as a point-type prompt for SAM2’s mask decoder, which then generates the final object mask.
@@ 3.2 Handling Tracking Loss in FoundationPose
FoundationPose relies on positional information from the previous frame for pose estimation, which can result in tracking loss if the object moves too quickly. To address this, SAM2 performs real-time object segmentation for each frame.
Figure 3: The Process of Robust tracking The image illustrates the process of robust tracking. In the first step, the frame is simultaneously sent to FoundationPose to obtain the pose estimation and to SAM2 to generate a mask. Afterward, the robust Lorentzian centroid distance is calculated. If this distance exceeds a predefined threshold, the system re-registers to solve the problem of tracking loss.
We calculate the distance between the centroid of the mask and the center point of the pose estimate to detect tracking loss as demonstrated in Figure 3. In Eq.1, this distance is measured using Robust Lorentzian centroid Distance with Lorentzian error function and square of L2 norm.
@@ 3.3 Addressing Long-Term Object Loss: A Memory Mechanism
To handle long-term object loss, we implemented a memory-like mechanism demonstrated in Figure 4. As defined in Eq.3, if the object is lost for an extended period (typically 10 seconds), both FoundationPose and SAM2 may fail to track it.
Figure 4: The process of memory mechanism The image depicts the memory mechanism process when an object is lost in the frame for an extended period. This is determined by measuring the difference between the area of the maximum contour and the initial contour. If the object remains lost for a specified duration, the segmented image is reloaded into LightGlue, which then passes a prompt to SAM2 to generate a mask for FoundationPose.
@@
In such cases, we reload the segmented image of the object saved during the initial SAM2 run and perform feature point matching on the current frames using LightGlue. If the object reappears, a location prompt is generated, which is passed to SAM2 for re-segmentation. The segmentation result is then fed into FoundationPose for re-registration, allowing the system to resume tracking the object.